OpenStudy (ledah):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=6, b=2, c=

OpenStudy (ledah):

how would I find this?

OpenStudy (hoblos):

c^2 = a^2 + b^2 so here c^2 = 6^2 + 2^2 = 36+4 = 4 then c= sqrt(40) = sqrt(4x10) = 2sqrt(10)

OpenStudy (ledah):

That doesn't really make sense to me.

OpenStudy (ledah):

Oh it didn't make sense because you made it backwards. I see it

OpenStudy (hoblos):

lol... dont you know Pythagoras theorem!! it says: in a right triangle: hypotenuse^2 = (first side)^2 + (second side)^2 use it here

OpenStudy (ledah):